Water is poured into a vessel of the form of a straight circular cone. If the vessel is installed with the <<sharp>> end down then the distance from the water level to the base of the cone is 1 m. When the vessel was turned over, it turned out that the distance from the water level to the <<sharp>> end of the vessel is root(3, 13) * m . Find the height of the vessel; give your answer in meters, rounded to two decimals if needed. The volume of a cone can be found by the V = 1/3 * S * h formula , where S is the area of its base and h is its height.


1
Expert's answer
2021-11-15T16:36:50-0500

Let "R=" the base radius of the cone, "H=" the height of the cone.

If the vessel is installed with the "sharp" end down

"x=" the base radius of the "water"cone, "h=" the height of the "water"cone.


"\dfrac{x}{h}=\dfrac{R}{H}"

The volume of the water is


"V_{water}=\dfrac{1}{3}\pi x^2h"

Given "h=1m"


"V_{water}=\dfrac{1}{3}\pi (\dfrac{R}{H})^2(1m)^3"

When the vessel was turned over


"V_{water}=\dfrac{1}{3}\pi R^2H-\dfrac{1}{3}\pi (\dfrac{R}{H})^2(\sqrt[3]{13}m)^3"

Then


"V_{water}=\dfrac{1}{3}\pi (\dfrac{R}{H})^2(1m)^3"

"=\dfrac{1}{3}\pi R^2H-\dfrac{1}{3}\pi (\dfrac{R}{H})^2(\sqrt[3]{13}m)^3"

"1=H^3-13"

"H=\sqrt[3]{14}\ m"

"H=2.41\ m"